# Tiny-Align: Bridging Automatic Speech Recognition and Large Language Model on the Edge

### v2

@@ Abstract
The combination of Large Language Models (LLM) and Automatic Speech Recognition (ASR), when deployed on edge devices (called edge ASR-LLM), can serve as a powerful personalized assistant to enable audio-based interaction for users. Compared to text-based interaction, edge ASR-LLM allows accessible and natural audio interactions. Unfortunately, existing ASR-LLM models are mainly trained in high-performance computing environments and produce substantial model weights, making them difficult to deploy on edge devices. More importantly, to better serve users’ personalized needs, the ASR-LLM must be able to learn from each distinct user, given that audio input often contains highly personalized characteristics that necessitate personalized on-device training. Since individually fine-tuning the ASR or LLM often leads to suboptimal results due to modality-specific limitations, end-to-end training ensures seamless integration of audio features and language understanding (cross-modal alignment), ultimately enabling a more personalized and efficient adaptation on edge devices. However, due to the complex training requirements and substantial computational demands of existing approaches, cross-modal alignment between ASR audio and LLM can be challenging on edge devices. In this work, we propose a resource-efficient cross-modal alignment framework that bridges ASR and LLMs on edge devices to handle personalized audio input. Our framework enables efficient ASR-LLM alignment on resource-constrained devices like NVIDIA Jetson Orin (8GB RAM), achieving 50x training time speedup while improving the alignment quality by more than 50%. To the best of our knowledge, this is the first work to study efficient ASR-LLM alignment on resource-constrained edge devices.
@@ I Introduction
The Large Language Model (LLM) deployed on the edge device (edge LLM) can serve as a personalized assistant, guaranteeing data privacy [1] and continuous service without relying on stable internet connections [2, 3]. Edge LLMs normally interact with users based on text (text-based interaction). Yet, it is more natural and accessible for users to interact with edge LLMs, their personalized assistants, via audio [4]. To enable the audio-based interaction, the Automatic Speech Recognition (ASR) models can be employed [5] to simply take the audio as the input and send the converted text into the edge LLM [6]. Existing works, on the other hand, have shown that this simple combination of ASR and LLM can fail in cases where the audio has no corresponding text [7], or the mismatching pre-trained knowledge in the ASR and the edge LLM can compromise the performance of their combination.
@@
Figure 1: Evaluation of ASR-LLM alignment using Whisper-base and Gemma-2-2B with and without GPTQ (G) compression. Comparison includes three existing approaches (A1, A2, A3) and a preliminary design using only projector and LLM embedding (Proj_only). This preliminary result motivates our work (detailed in Fig. 2), which outperforms existing approaches.
@@
For approaches A1 and A2, the LLM generation involved can consume a huge amount of resources, leading to a longer convergence time. The approach of A2, which takes the projector and the LLM for training, outperforms the other two approaches but it also takes the most resources. The approach of A1 is more lightweight, but it still uses the LLM, and its reduction in finetuning resources leads to significant performance degradation. Meanwhile, The A1’s lower performance than the A2 can be due to the simple projector design. As shown in Fig. 1, while the A2 obtains decent performance, it also has the highest latency mostly due to the LLM generation. The Proj_only is tentatively made to avoid the LLM generation. To do that, we pick and use only its embedding layer to avoid the LLM generation. The embedding layer interfaces directly with the ASR encoder through a well-designed projector, which requires both sufficient embedding capacity and architectural compatibility with the ASR encoder and the rest components of the edge LLM. In Fig. 1, we compare this design as Proj_only with the three existing approaches. Surprisingly, we discovered that this design can outperform A1, A2, and even A3 in terms of performance and training latency.
@@ II-A Automatic Speech Recognition onto edge LLM
Wav2vec2 [15] pioneered an efficient approach to audio processing by directly learning representations from raw waveforms. Through a multi-layer convolutional network, it extracts context-dependent representations that effectively capture speech patterns. The model employs contrastive learning to identify relevant features and discern patterns in the audio signal. While its streamlined architecture makes it particularly efficient for edge deployment, wav2vec2 may struggle with highly noisy environments.
@@ II-B Cross-modal Alignment on Edge
One key challenge in personalizing edge LLM is adapting to users’ preferred input modality while operating within the constraints of limited computational resources [21, 22, 23]. Cross-modal alignment enables the integration of non-textual inputs by combining specialized machine learning models for various modalities with the edge LLM which provides advanced reasoning capabilities for analyzing cross-modal data [24]. However, this alignment process demands substantial resources, making deployment on edge devices particularly challenging [25, 26]. For example, on resource-constrained devices like NVIDIA Jetson Orin (275 TOPS, INT8), cross-modal alignment can quickly exhaust available memory and computational resources, with processing latency becoming prohibitive for real-time applications. These constraints highlight the need for optimizing cross-modal alignment specifically for edge deployment.
@@ IV-A2 Default Experimental Setting
The training process uses the AdamW optimizer with an initial learning rate of 1e-3 and a linear decay schedule. This configuration balances training stability with convergence speed, particularly important for edge deployment scenarios where training resources may be limited. The experiments are run on a single Nvidia P100 GPU, whose memory and computational power are close to edge devices like Jetson Orin AGX or Dell Precision 7560 with Laptop GPU.
IV-A3 Evaluation Methods
We evaluate our framework from both effectiveness and efficiency perspectives. For effectiveness, we employ a dual-path comparison approach: given an audio input, we compare the output from our ASR-LLM system (OutputA) with the output generated by directly feeding the ground truth transcript to the LLM (OutputL). The similarity between these outputs is measured using ROUGE scores [41], including ROUGE-1 and ROUGE-L, which assess word-level and longest common subsequence matches respectively. For the ROUGE scores, the higher values of them represent the better performance of the LLM generated content.
For efficiency evaluation, we focus on convergence time as our primary metric, which measures not only the training speed but also resource utilization efficiency [42]. Extended training periods on edge devices can lead to sustained high DRAM usage, GPU occupancy, and power consumption. We define convergence as the point where loss stabilizes ($\Delta\text{loss}<\varepsilon$), analyzing both time-to-convergence and total resource consumption during training. This metric is particularly crucial for edge deployment, where prolonged training can strain limited device resources and impact system availability. To evaluate personalization effectiveness, we conduct user-level performance analysis by randomly sampling 20 users from each dataset and computing their average performance metrics.
IV-A4 Baselines
@@
IV-B Results
In our experiments, we first analyze the training loss convergence of our method compared to three baselines. We evaluate an ASR-LLM system composed of wav2vec2 and Gemma-2-2B on two datasets: ADReSS-IS2020 and ENNI. Training loss is tracked over 1000 epochs. As shown in Fig. 6, our method achieves convergence within 400 epochs on ADReSS-IS2020 and demonstrates even faster convergence (within 100 epochs) on the ENNI dataset. This rapid convergence enables efficient ASR-LLM alignment while minimizing computational resource usage and training time. Additionally, we can observe the drastic fluctuations of the loss curve for the three baselines, indicating the training under these three methods can be less stable. One thing to notice is that, since the baselines are designed without considering the resource restrictions, we must limit their trainable parameters so they can run in resource-limited environments.
Following our analysis of data size impact, we conduct a comprehensive evaluation of our method against the baselines. Based on our training loss analysis, we set the training epoch limit to 400 to reflect resource-constrained scenarios in ASR-LLM alignment. Given that our previous study showed no significant performance variations across different data sizes, we use a dataset of 100 samples for this comprehensive evaluation. As shown in TABLE I, our proposed method, Tiny-Align, demonstrates superior performance across five datasets and five ASR-LLM systems, achieving significantly higher ROUGE-1 and ROUGE-L scores compared to the three baselines. Furthermore, we analyze convergence time without the 400-epoch constraint, setting a target loss of 0.05. The results, presented in the third column for each ASR-LLM system in TABLE I, show that Tiny-Align achieves significantly faster convergence compared to the baseline methods.
@@ IV-B Results
Additionally, we evaluate the instruction injection mechanism on two ASR-LLM combinations: wav2vec2+Gemma-2-2B and wav2vec2+StableLM-2-1.6B. In these experiments, we remove the epoch limit to allow complete BridgeFormer training. Instructions are then injected during the inference stage. As shown in Fig. 8, instruction injection significantly improves LLM generation performance with the same trained BridgeFormer, indicating enhanced LLM comprehension of BridgeFormer-produced audio embeddings.

### v3

# Tiny-Align: Bridging Automatic Speech Recognition and Large Language Model on Edge

@@ Abstract
The combination of Large Language Models (LLM) and Automatic Speech Recognition (ASR), when deployed on edge devices (called edge ASR-LLM), can serve as a powerful personalized assistant to enable audio-based interaction for users. Compared to text-based interaction, edge ASR-LLM allows accessible and natural audio interactions. Unfortunately, existing ASR-LLM models are mainly trained in high-performance computing environments and produce substantial model weights, making them difficult to deploy on edge devices. More importantly, to better serve users’ personalized needs, the ASR-LLM must be able to learn from each distinct user, given that audio input often contains highly personalized characteristics that necessitate personalized on-device training. Since individually fine-tuning the ASR or LLM often leads to suboptimal results due to modality-specific limitations, end-to-end training ensures seamless integration of audio features and language understanding (cross-modal alignment), ultimately enabling a more personalized and efficient adaptation on edge devices. However, due to the complex training requirements and substantial computational demands of existing approaches, cross-modal alignment between ASR audio and LLM can be challenging on edge devices. In this work, we propose a resource-efficient cross-modal alignment framework that bridges ASR and LLMs on edge devices to handle personalized audio input. Our framework enables efficient ASR-LLM alignment on resource-constrained devices like Raspberry Pi 5 (8GB RAM), achieving 50x training time speedup while improving the alignment quality by more than 50%. To the best of our knowledge, this is the first work to study efficient ASR-LLM alignment on resource-constrained edge devices.
@@ I Introduction
The Large Language Model (LLM) deployed on the edge device (edge LLM) can serve as a personalized assistant, guaranteeing data privacy [1] and continuous service without relying on stable internet connections [2, 3, 4]. Edge LLMs normally interact with users based on text (text-based interaction). Yet, it is more natural and accessible for users to interact with edge LLMs, their personalized assistants, via audio [5]. To enable the audio-based interaction, the Automatic Speech Recognition (ASR) models can be employed [6] to simply take the audio as the input and send the converted text into the edge LLM [7]. Existing works, on the other hand, have shown that this simple combination of ASR and LLM can fail in cases where the audio has no corresponding text [8], or the mismatching pre-trained knowledge in the ASR and the edge LLM can compromise the performance of their combination.
Figure 1: Evaluation of ASR-LLM alignment using Whisper-base and Gemma-2-2B with and without GPTQ (G) compression. Comparison includes three existing approaches (A1, A2, A3) and a preliminary design using only projector and LLM embedding (off-the-Shelf). This preliminary result motivates our work (detailed in Fig. 2), which outperforms existing approaches.
@@
For approaches A1 and A2, the LLM generation involved can consume a huge amount of resources, leading to a longer convergence time. The approach of A2, which takes the projector and the LLM for training, outperforms the other two approaches but it also takes the most resources. The approach of A1 is more lightweight, but it still uses the LLM, and its reduction in finetuning resources leads to significant performance degradation. Meanwhile, The A1’s lower performance than the A2 can be due to the simple projector design. As shown in Fig. 1, while the A2 obtains decent performance, it also has the highest latency mostly due to the LLM generation. The off-the-shelf is tentatively made to avoid the LLM generation. To do that, we pick and use only its embedding layer to avoid the LLM generation. The embedding layer interfaces directly with the ASR encoder through a well-designed projector, which requires both sufficient embedding capacity and architectural compatibility with the ASR encoder and the rest components of the edge LLM. In Fig. 1, we compare this design as off-the-shelf with the three existing approaches. Surprisingly, we discovered that this design can outperform A1, A2, and even A3 in terms of performance and training latency.
@@ II-A Automatic Speech Recognition onto edge LLM
Wav2vec2 [16] pioneered an efficient approach to audio processing by directly learning representations from raw waveforms. Through a multi-layer convolutional network, it extracts context-dependent representations that effectively capture speech patterns. The model employs contrastive learning to identify relevant features and discern patterns in the audio signal. While its streamlined architecture makes it particularly efficient for edge deployment, wav2vec2 has demonstrated robust performance in various scenarios [17], including child speech recognition, highlighting its capability to effectively capture speech features across diverse conditions.
@@ II-B Cross-modal Alignment on Edge
One key challenge in personalizing edge LLM is adapting to users’ preferred input modality while operating within the constraints of limited computational resources [23, 24, 25, 26]. Cross-modal alignment enables the integration of non-textual inputs by combining specialized machine learning models for various modalities with the edge LLM which provides advanced reasoning capabilities for analyzing cross-modal data [27]. However, this alignment process demands substantial resources, making deployment on edge devices particularly challenging [28, 29]. For example, on resource-constrained devices like Raspberry Pi 5 (8GB RAM, quad-core Cortex-A76), cross-modal alignment can quickly exhaust available memory and computational resources, with processing latency becoming prohibitive for real-time applications. These constraints highlight the need for optimizing cross-modal alignment specifically for edge deployment.
@@ IV-A3 Default Experimental Setting
The training process uses the AdamW optimizer with an initial learning rate of 1e-3 and a linear decay schedule. This configuration balances training stability with convergence speed, which is particularly important for edge deployment scenarios where training resources may be limited.
The preliminary validation on a Rasberry Pi 5 can be found in section IV-B1, followed by the comprehensive experiments run on a single Nvidia P100 GPU.
IV-A4 Evaluation Methods
We evaluate our framework from both effectiveness and efficiency perspective. For effectiveness, we employ a dual-path comparison approach: given an audio input, we compare the output from our ASR-LLM system (OutputA) with the output generated by directly feeding the ground truth transcript to the LLM (OutputL). The similarity between these outputs is measured using ROUGE scores [44], including ROUGE-1 and ROUGE-L, which assess word-level and longest common subsequence matches, respectively. For the ROUGE scores, higher values represent better performance of the LLM generated content.
For efficiency evaluation, we focus on convergence time as our primary metric, which measures not only the training speed but also resource utilization efficiency [45]. Extended training periods on edge devices can lead to sustained high DRAM usage, GPU occupancy, and power consumption. We define convergence as the point where the loss stabilizes ($\Delta\text{loss}<\varepsilon$), analyzing both time-to-convergence and total resource consumption during training. This metric is particularly crucial for edge deployment, where prolonged training can strain limited device resources and impact system availability. To evaluate personalization effectiveness, we conduct user-level performance analysis by randomly sampling 20 users from each dataset and computing their average performance metrics.
IV-A5 Baselines
@@
Figure 6: Demonstration of the viability of Tiny-Align on a resource-constrained edge device (Raspberry Pi 5), evaluated under various configurations (x-axis) of the proposed BridgeFormer. The default settings are marked in dark green.
IV-B Results
IV-B1 Edge Device Implementation Validation
To assess whether edge device can support running the full model pipeline (i.e., ASR, projector, and LLM), we perform a preliminary validation on a Raspberry Pi 5. In this experiment, we evaluate two critical factors: throughput and RAM usage.
As shown in Figure 6, we explore various configurations of our BridgeFormer. The left panel illustrates the trends in throughput and memory usage across different hidden dimensions, while the right panel shows the impact of varying the number of attention heads.
IV-B2 Comprehensive Experiments
In this set of the experiments, we first analyze the training loss convergence of our method compared to three baselines. We evaluate an ASR-LLM system composed of wav2vec2 and Gemma-2-2B on two datasets: ADReSS-IS2020 and ENNI. Training loss is tracked over 1000 epochs. As shown in Fig. 7, our method achieves convergence within 400 epochs on ADReSS-IS2020 and demonstrates even faster convergence (within 100 epochs) on the ENNI dataset. This rapid convergence enables efficient ASR-LLM alignment while minimizing computational resource usage and training time. Additionally, we can observe the drastic fluctuations of the loss curve for the three baselines, indicating the training under these three methods may be less stable. One thing to note is that, since the baselines are designed without considering the resource restrictions, we must limit their trainable parameters so that they can run in resource-limited environments.
@@ IV-B2 Comprehensive Experiments
Following our analysis of data size impact, we conduct a comprehensive evaluation of our method against the baselines. Based on our training loss analysis, we set the training epoch limit to 400 to reflect resource-constrained scenarios in ASR-LLM alignment. Given that our previous study did not show significant performance variations for different data sizes, we use a dataset of 100 samples for this comprehensive evaluation. As shown in TABLE I, our proposed method, Tiny-Align, demonstrates superior performance across five datasets and five ASR-LLM systems, achieving significantly higher ROUGE-1 and ROUGE-L scores compared to the three baselines. Furthermore, we analyze convergence time without the 400-epoch constraint, setting a target loss of 0.05. The results, presented in the third column for each ASR-LLM system in TABLE I, show that Tiny-Align achieves significantly faster convergence compared to the baseline methods.
@@
Furthermore, we evaluate the instruction injection mechanism on two ASR-LLM combinations: wav2vec2+Gemma-2-2B and wav2vec2+StableLM-2-1.6B. In these experiments, we remove the epoch limit to allow complete BridgeFormer training. Instructions are then injected during the inference stage. As shown in Fig. 9, instruction injection significantly improves LLM generation performance with the same trained BridgeFormer, indicating enhanced LLM comprehension of BridgeFormer-produced audio embeddings.